OpenStudy (anonymous):

Solve for x. x2 - 7x = 0 A. 0, -7 B. 0, 7 C. 1, -7 D. 1, 7

OpenStudy (anonymous):

B. 0,7

OpenStudy (anonymous):

it shows you the graph :D

OpenStudy (zzr0ck3r):

x^2 - 7x = 0 factor out an x x(x-7) = 0 x = 0 and x-7 = 0 so x = 0, x = 7

OpenStudy (anonymous):

thanks :)